OpenStudy (anonymous):

Place the indicated product in the proper location on the grid. (3 - c2d)(4 - 4c2d)

OpenStudy (anonymous):

\[\left(4-4 c^2 d\right) \left(3-c^2 d\right)=4 c^4 d^2-16 c^2 d+12 \]

OpenStudy (anonymous):

It's a shame this answer is incorrect.